jovanie:

a square rug covers 79 square feet of floor. what is the approximate length of one side of the rug. nearest hundreth foot

jovanie:

please help me please

Mercury:

old question but will respond to help out future users if you are given the area of a square, and you want to find the length of one side, simply take the square root of the area A = s^2 therefore: s = sqrt(A)